# Preference Learning Algorithms Do Not Learn Preference Rankings

### v1

@@ Abstract
Preference learning algorithms (e.g., RLHF and DPO) are frequently used to steer LLMs to produce generations that are more preferred by humans, but our understanding of their inner workings is still limited. In this work, we study the conventional wisdom that preference learning trains models to assign higher likelihoods to more preferred outputs than less preferred outputs, measured via ranking accuracy. Surprisingly, we find that most state-of-the-art preference-tuned models achieve a ranking accuracy of less than $60\%$ on common preference datasets. We furthermore derive the idealized ranking accuracy that a preference-tuned LLM would achieve if it optimized the DPO or RLHF objective perfectly. We demonstrate that existing models exhibit a significant alignment gapi.e., a gap between the observed and idealized ranking accuracies. We attribute this discrepancy to the DPO objective, which is empirically and theoretically ill-suited to fix even mild ranking errors in the reference model, and derive a simple and efficient formula for quantifying the difficulty of learning a given preference datapoint. Finally, we demonstrate that ranking accuracy strongly correlates with the empirically popular win rate metric when the model is close to the reference model used in the objective, shedding further light on the differences between on-policy (e.g., RLHF) and off-policy (e.g., DPO) preference learning algorithms.
@@ 1 Introduction
    Existing models do not achieve high ranking accuracies. We demonstrate that a wide variety of open-access preference-tuned LLMs (e.g., Llama 2 7B Chat, Gemma 7B IT, and Zephyr 7B DPO) achieve a ranking accuracy below $60\%$ across a range of validation splits from commonly used preference datasets, such as UltraFeedback [6], Anthropic helpfulness and harmlessness (HH-RLHF, [13]), and Stanford Human Preferences (SHP, [10]) (Figure 1).
@@ 2.1 Learning from Human Preferences
${\mathcal{L}}_{\text{DPO}}({\pi_{\theta}},{\pi_{\text{Ref}}})$ $=-\mathop{\mathbb{E}}_{(x,y_{w},y_{l})\sim{\mathcal{D}}}[\log\sigma(\underbrace{\beta\log\frac{{\pi_{\theta}}(y_{w}\mid x)}{{\pi_{\text{Ref}}}(y_{w}\mid x)}-\beta\log\frac{{\pi_{\theta}}(y_{l}\mid x)}{{\pi_{\text{Ref}}}(y_{l}\mid x)}}_{\text{reward margin}})]$
$=-\mathop{\mathbb{E}}_{(x,y_{w},y_{l})\sim{\mathcal{D}}}[\log\sigma(\beta\underbrace{\log\frac{{\pi_{\theta}}(y_{w}\mid x)}{{\pi_{\theta}}(y_{l}\mid x)}}_{\text{model log-ratio}}+\beta\underbrace{\log\frac{{\pi_{\text{Ref}}}(y_{l}\mid x)}{{\pi_{\text{Ref}}}(y_{w}\mid x)}}_{\text{reference modellog-ratio}})]$
@@ 2.2 Evaluation Metrics
Analogously, the ranking accuracy of policy $\pi_{\theta}$ on a dataset $\mathcal{D}=\{(x,y_{w},y_{l})\}$ is ${{\mathcal{R}}}(\mathcal{D};\pi_{\theta})={\mathop{\mathbb{E}}}_{(x,y_{w},y_{l})\sim\mathcal{D}}{{\mathcal{R}}}(x,y_{w},y_{l};\pi_{\theta}).$ In the rare case where a dataset has more than two outputs $y$ per prompt $x$, we use the generalized ranking accuracy definition stated in App. A.6.
@@ 3.1 Existing Reference Models Rarely Have Correct Rankings
Figure 1: Both reference and preference-tuned models exhibit low ranking accuracy on most preference datasets. Each point represents the length-normalized or non-length-normalized ranking accuracy of individual (1(a)) reference models (pre-trained or fine-tuned), or (1(b)) preference-tuned models (trained with DPO or RLHF). We sub-sample 1K examples from each dataset and use the test split when available. We describe datasets in C.2 and list all numbers in Tables 2, 3, and 4.
Fig. 1(a) indicates that reference models rarely achieve high ranking accuracy on common preference datasets (except Synthetic Instruct GPT-J Pairwise), even though many were likely trained on the preferred completions (see Sef. 2.1). Many of the models do not have documented training data so we do not know which preference datasets, if any, are in-distribution. We also fine-tune several pretrained LLMs on the preferred completions (see App. C.1) and observe that ranking accuracy does not increase significantly. [^2] It is not surprising that fine-tuning on the preferred completions does not boost ranking accuracy, since the model does not receive any knowledge of the relative qualities of the preferred and rejected completions. Based on our findings, we turn to the case of imperfect reference models.
@@ 3.2 Idealized Ranking Accuracy
We prove this result in App. A.2. This theorem highlights the importance of reporting multiple ratings for each preference datapoint, because it shows that when $\alpha(x,y_{w},y_{l})=1$, the objectives only have an optimum at infinity.
@@ 3.3 Measuring the Alignment Gap
Given access to $\pi_{\text{Ref}}$, $\beta$, and the $\alpha(x,y_{w},y_{l})$ values for each triple $(x,y_{w},y_{l})$ in a given preference dataset, we can compute the idealized ranking accuracy from Eq. 3.555Note that when $\alpha(x,y_{w},y_{l})=1$, we replace it with $1-\epsilon$ to compute the formula. See App. C.4. The results are shown in Table 1 and further details are given in App. C.4.
@@
Table 1: The idealized ranking accuracy of existing algorithms is not perfect, but preference-tuned models exhibit ranking accuracies far even from this idealized case. We provide both the length-normalized ($\tilde{{\mathcal{R}}}$) and non-length-normalized (${{\mathcal{R}}}$) ranking accuracies for a variety of open-access preference-tuned models on the Alpaca Farm [8] validation dataset (described in App. C.2). We also provide the idealized ranking accuracy (Corollary 3.3). Since idealized ranking accuracy can be computed with a variety of values of $\beta$, we provide the minimum, median, and maximum idealized ranking accuracy values for a range of $\beta$. For more details, see App. C.4.
@@ 4.1 DPO Rarely Flips Preference Rankings
Figure 2: Despite continuously decreasing the loss, DPO rarely flips the rankings of pairs and instead mostly increases the reward margin of already correctly ranked pairs. We train a Pythia-2.8B model using the DPO objective and categorize the training dataset into four subsets – examples that initially have the correct ranking and are flipped to (1) correct or (2) incorrect, and examples that initially have the incorrect ranking and are flipped to (3) correct or (4) incorrect. In all three figures, the hue of the point indicates the category. The dashed vertical line indicates the training step at which the lowest eval. loss occurs. We also present results for two other models with three seeds each in Appendix D.
To study how ranking accuracy changes over the course of DPO training, we train three sizes of models (GPT-2 [38], Pythia 2.8B [4], and Llama 2-7B [50]) across three seeds each on the Anthropic HH-RLHF [3] preference dataset and study the ranking accuracy on different partitions of the training dataset. We present the results from training one seed of Pythia 2.8B in Fig. 2, and defer training details to App. D.1 and results on the other two models to App. D.2. In Fig. 2, we partition a random subsample of 1K examples from the training dataset into four groups based on whether the reference model ${\pi_{\text{Ref}}}$ had the correct ranking and whether the current model ${\pi_{\theta}}$ has the correct ranking.
Surprisingly, Fig. 2 demonstrates that DPO rarely flips the ranking of $(y_{w},y_{l})$ over the course of training despite consistently reducing the loss ${\mathcal{L}}_{\text{DPO}}$. Aside from the group of points for which the model unlearns the correct preference ranking, we observe that the loss decreases and the reward margin increases consistently while training. However, at the point of lowest validation loss, less than half of the incorrectly ranked points have been flipped to have the correct ranking. This indicates that the DPO objective is ill-formulated to induce a high ranking accuracy in practice.
@@ 4.2 Analysis: How Easy Is It To Flip A Ranking?
We prove this result in App. A.4. Our theoretical result allows us to formally identify the points that will be hard to flip in their rankings. Fig. 3 visualizes the reference model log-ratio for several settings and highlights that datapoints with even mild ranking errors in the reference model will require the loss to be reduced to a very low value in order to flip the ranking. App. F contains examples of hard-to-learn, easy-to-learn, and easy-to-flip datapoints. We observe that the hard-to-learn datapoints are substantially longer than the easy ones, and that the easy datapoints generally contain less ambiguous preference annotations. More generally, our result motivates the use of stronger $\pi_{\text{Ref}}$ models and iterative or on-policy variants of DPO [48, 63, 23, 51].
@@ 5 Ranking Accuracy and Win Rate
We speculate that when the model is far from the reference model, regularization toward the reference model can harm the generative capabilities of the model, which are primarily acquired during pre-training. In other words, the off-policy behavior of the model can no longer predictably describe the on-policy generations when the reference model used in the offline objective is far from the current model. Our findings confirm the fundamental tradeoff between fitting the preference data and maintaining generative capabilities acquired during pre-training [21] and align with prior observations that adding on-policy preference data can make offline learning more effective [48, 63, 23, 51].
(a) Ranking accuracy and win rate of various Pythia 2.8B checkpoints during training, versus the distance travelled by the model weights $\theta_{t}$ from the initialization.
(b) Ranking accuracy and win rate of various $\gamma$-scaled models, versus the distance travelled by the model weights $\theta_{\gamma}$ from the initialization.
Figure 4: When the model weights have not travelled far from $\theta_{\text{Ref}}$, ranking accuracy and win rate increase together. $\theta_{t}$ represents the model weights at checkpoint $t$, and $\theta_{\gamma}$ are the model weights for the model trained with ${\mathcal{L}}_{\text{DPO}}^{\gamma}$.
@@ 6 Related Work
Many works have investigated the role of the preference dataset [54, 59], the reliability of the evaluations [67, 28], and the confounding factor of response length [43, 9, 55, 35]. Theoretical works have unified the many preference learning algorithms into clear taxonomies that permit analysis and, sometimes, yield new variants [2, 59, 46, 48, 61]. Several works study idiosyncrasies of preference learning, such as why DPO decreases the likelihood of both rejected and chosen outputs from the dataset [40, 12, 33] and why RLHF exhibits vanishing gradients [42]. In contrast, our work approaches understanding DPO and RLHF through the lens of ranking accuracy, and our findings emphasize the role of the reference model regularization in preference learning. Relatedly, SliC-HF [65], CPO [57], and pairwise cringe loss [58] optimize log probability margins $\log\pi(y_{w}|x)-\log\pi(y_{l}|x)$, effectively removing the regularization toward the reference model. Liu et al. [31] recommend using the reference model at inference time to exert more granular control over the regularization. Tang et al. [48] also analyzed the role of regularizing toward a reference model, though our work focuses the effect of this regularization on ranking accuracy. App. B discusses additional related works.
@@ 7 Discussion
Our work highlights the significant but nuanced relationship between preference learning and ranking accuracy. We have demonstrated both theoretically and empirically that RLHF and DPO struggle to teach the model to correctly rank preferred and dispreferred outputs, even in the training dataset. Although the learning objective promotes high ranking accuracy in theory (Proposition 2.6), we observed a prominent alignment gap resulting from the poor conditioning of reference models. We then drew connections between the off-policy nature of ranking accuracy and the on-policy evaluations of win rate, identifying specific scenarios in which on-policy behavior can or cannot be reliably predicted by off-policy behavior. App. B details the limitations of our work.
@@ 8 Acknowledgements
We thank Sanjeev Arora, Tianyu Gao, Eric Mitchell, Richard Pang, and Mengzhou Xia for helpful discussions during the development of this project. We also thank Nikita Nangia, Eshaan Nichani, and Alexander Wettig for their help in proofreading the work.
This work was supported by National Science Foundation Award 1922658, the Samsung Advanced Institute of Technology (under the project Next Generation Deep Learning: From Pattern Recognition to AI), NSF CAREER Award 2145542, ONR N00014-23-1-2634, and Google. SM is additionally supported by ONR, NSF, and DARPA. We also thank Princeton Language and Intelligence (PLI) for computing resources and OpenAI credits, and NYU High Performance Computing (HPC) for computing resources and in-kind support.

### v4

@@ Abstract
Preference learning algorithms (e.g., RLHF and DPO) are frequently used to steer LLMs to produce generations that are more preferred by humans, but our understanding of their inner workings is still limited. In this work, we study the conventional wisdom that preference learning trains models to assign higher likelihoods to more preferred outputs than less preferred outputs, measured via ranking accuracy. Surprisingly, we find that most state-of-the-art preference-tuned models achieve a ranking accuracy of less than $60\%$ on common preference datasets. We also derive the idealized ranking accuracy that a preference-tuned LLM would achieve if it optimized the DPO or RLHF objective perfectly. We demonstrate that existing models exhibit a significant alignment gapi.e., a gap between the observed and idealized ranking accuracies. We attribute this discrepancy to the DPO objective, which is empirically and theoretically ill-suited to fix even mild ranking errors in the reference model, and derive a simple and efficient formula for quantifying the difficulty of learning a given preference datapoint. Finally, we show that ranking accuracy strongly correlates with the empirically popular win rate metric when the model is close to the reference model, shedding further light on the differences between on-policy (e.g., RLHF) and off-policy (e.g., DPO) preference learning algorithms.
@@ 1 Introduction
    Existing models do not achieve high ranking accuracies. We demonstrate that a wide variety of open-access preference-tuned LLMs (e.g., Llama 2 7B Chat, Gemma 7B IT, and Zephyr 7B DPO) achieve a ranking accuracy below $60\%$ across a range of validation splits from commonly used preference datasets, such as UltraFeedback [7], Anthropic helpfulness and harmlessness (HH-RLHF, [14]), and Stanford Human Preferences (SHP, [11]) (Figure 1). Although we do not advocate for ranking accuracy as a measure of model quality, we analyze LLMs’ ranking accuracies nonetheless because (1) ranking accuracy has motivated the design of many preference learning algorithms, and (2) DPO directly optimizes for ranking accuracy (Theorem 3.1).
@@ 2.1 Learning from Human Preferences
${\mathcal{L}}_{\text{DPO}}({\pi_{\theta}},{\pi_{\text{Ref}}})$ $=-\mathop{\mathbb{E}}_{(x,y_{w},y_{l})\sim{\mathcal{D}}}\biggl{[}\log\sigma\biggl{(}\underbrace{\beta\log\frac{{\pi_{\theta}}(y_{w}\mid x)}{{\pi_{\text{Ref}}}(y_{w}\mid x)}-\beta\log\frac{{\pi_{\theta}}(y_{l}\mid x)}{{\pi_{\text{Ref}}}(y_{l}\mid x)}}_{\text{reward margin}}\biggr{)}\biggr{]}$
$=-\mathop{\mathbb{E}}_{(x,y_{w},y_{l})\sim{\mathcal{D}}}\biggl{[}\log\sigma\biggl{(}\beta\underbrace{\log\frac{{\pi_{\theta}}(y_{w}\mid x)}{{\pi_{\theta}}(y_{l}\mid x)}}_{\text{model log-ratio}}+\beta\underbrace{\log\frac{{\pi_{\text{Ref}}}(y_{l}\mid x)}{{\pi_{\text{Ref}}}(y_{w}\mid x)}}_{\text{reference model log-ratio}}\biggr{)}\biggr{]}$
@@ 2.2 Evaluation Metrics
Analogously, the ranking accuracy of policy $\pi_{\theta}$ on a dataset $\mathcal{D}=\{(x,y_{w},y_{l})\}$ is ${{\mathcal{R}}}(\mathcal{D};\pi_{\theta})={\mathop{\mathbb{E}}}_{(x,y_{w},y_{l})\sim\mathcal{D}}{{\mathcal{R}}}(x,y_{w},y_{l};\pi_{\theta}).$ In the rare case where a dataset has more than two outputs $y$ per prompt $x$, we use the generalized ranking accuracy definition stated in App. A.6. We do not advocate for ranking accuracy as a metric of model quality, but as a lens into the inner workings of common preference learning algorithms.
@@ 3.1 Existing Reference Models Rarely Have Correct Rankings
Figure 1: Both reference and preference-tuned models exhibit low ranking accuracy on most preference datasets. Each point represents the length-normalized or non-length-normalized ranking accuracy of individual (1(a)) reference models (pre-trained or fine-tuned), or (1(b)) preference-tuned models (trained with DPO or RLHF). The random chance accuracy for each dataset is indicated with a black ‘X’. We sub-sample 1K examples from each dataset and use the test split when available. We describe datasets in B.2 and list all numbers in Tables 2, 3, and 4. For UltraFeedback, ranking accuracy is measured with exact match across all 4 outputs (see App. A.6).
Fig. 1(a) indicates that reference models rarely achieve high ranking accuracy on common preference datasets (except Synthetic Instruct GPT-J Pairwise), even though many were likely trained on the preferred completions (see Sef. 2.1). Many of the models do not have documented training data so we do not know which preference datasets, if any, are in-distribution. We also fine-tune several pretrained LLMs on the preferred completions (see App. B.1) and observe that ranking accuracy does not increase significantly. [^2] It is not surprising that fine-tuning on the preferred completions does not boost ranking accuracy, since the model does not receive any knowledge of the relative qualities of the preferred and rejected completions. Based on our findings, we turn to the case of imperfect reference models.
@@ 3.2 Idealized Ranking Accuracy
We prove this result in App. A.2. Note that deterministic preferences, i.e., $\alpha(x,y_{w},y_{l})=1$, should not be confused with settings that only report a single individual’s preference, i.e., $\hat{\alpha}(x,y_{w},y_{l})=1$. In the former, the optimal probability ratio is infinity. The latter requires a better estimate of $\mathbb{P}(y_{w}\succ y_{l})$, e.g., more samples.
@@ 3.3 Measuring the Alignment Gap
Given access to $\pi_{\text{Ref}}$, $\beta$, and the $\alpha(x,y_{w},y_{l})$ values for each triple $(x,y_{w},y_{l})$ in a given preference dataset, we can compute the idealized ranking accuracy from Eq. 3.555Note that when $\alpha(x,y_{w},y_{l})=1$, we replace it with $1-\epsilon$ to compute the formula. See App. B.4. The results are shown in Table 1 and further details are given in App. B.4.
@@
Table 1: The idealized ranking accuracy of existing algorithms is not perfect, but preference-tuned models exhibit ranking accuracies far even from this idealized case. We provide both the length-normalized ($\tilde{{\mathcal{R}}}$) and non-length-normalized (${{\mathcal{R}}}$) ranking accuracies for a variety of open-access preference-tuned models on the Alpaca Farm [9] validation dataset (described in App. B.2). We also provide the idealized ranking accuracy (${{\mathcal{R}}^{*}}$ or $\tilde{{\mathcal{R}}^{*}}$, Corollary 3.3). Since idealized ranking accuracy can be computed with a variety of values of $\beta$, we provide the minimum, median, and maximum idealized ranking accuracy values for a range of $\beta$. For more details, see App. B.4.
@@ 4.1 DPO Rarely Flips Preference Rankings
Figure 2: Despite continuously decreasing the loss, DPO rarely flips the rankings of pairs before the point of overfitting (marked by the vertical dashed line) and instead mostly increases the reward margin of already correctly ranked pairs. We train a Pythia-2.8B model for 5 epochs using the DPO objective and categorize the training dataset into four subsets – examples that initially have the correct ranking and are flipped to (1) correct or (2) incorrect, and examples that initially have the incorrect ranking and are flipped to (3) correct or (4) incorrect. In all three figures, the hue of the point indicates the category. The dashed vertical line indicates the training step at which the lowest eval. loss occurs. Past this point, the model begins to overfit (i.e., the eval. loss starts to increase). We also present results for two other models with three seeds each in Appendix C.
To study how ranking accuracy changes over the course of DPO training, we train three sizes of models (GPT-2 [40], Pythia 2.8B [4], and Llama 2-7B [52]) across three seeds each on the Anthropic HH-RLHF [3] preference dataset and study the ranking accuracy on different partitions of the training dataset. We present the results from training one seed of Pythia 2.8B in Fig. 2, and defer training details to App. C.1 and results on the other two models to App. C.2. In Fig. 2, we partition a random subsample of 1K examples from the training dataset into four groups based on whether the reference model ${\pi_{\text{Ref}}}$ had the correct ranking and whether the current model ${\pi_{\theta}}$ has the correct ranking.
Surprisingly, Fig. 2 demonstrates that DPO rarely flips the ranking of $(y_{w},y_{l})$ over the course of training despite consistently reducing the loss ${\mathcal{L}}_{\text{DPO}}$. Aside from the group of points for which the model unlearns the correct preference ranking, we observe that the loss decreases and the reward margin increases consistently while training. However, at the point of lowest validation loss (marked by the vertical dashed line in Fig. 2(c)), less than $10\%$ of the originally incorrectly ranked points have been flipped to have the correct ranking. Past this point, the model begins to overfit (i.e., the validation loss begins to increase). DPO does not substantially improve the ranking accuracy until far past the point of overfitting. Although Theorem 3.1 predicts that the theoretically optimal DPO model exhibits high ranking accuracy, in practice the empirical endpoint of training occurs far from the theoretical optimum. This indicates that the DPO objective is ill-formulated to induce a high ranking accuracy in practice.
@@ 4.2 Analysis: How Easy Is It To Flip A Ranking?
We prove this result in App. A.4. Our theoretical result allows us to formally identify the points that will be hard to flip in their rankings. Fig. 3 visualizes the reference model log-ratio for several settings and highlights that datapoints with even mild ranking errors in the reference model will require the loss to be reduced to a very low value in order to flip the ranking. App. E contains examples of hard-to-learn, easy-to-learn, and easy-to-flip datapoints. We observe that the hard-to-learn datapoints are substantially longer than the easy ones, and that the easy datapoints generally contain less ambiguous preference annotations. More generally, our result motivates the use of stronger $\pi_{\text{Ref}}$ models and iterative or on-policy variants of DPO [50, 65, 24, 53].
@@ 5 Ranking Accuracy and Win Rate
We speculate that when the model is far from the reference model, overly optimizing the reward margin can harm the generative capabilities of the model, which are primarily acquired during pre-training. In other words, the off-policy behavior of the model can no longer predictably describe the on-policy generations when the reference model used in the offline objective is far from the current model. Our findings confirm the fundamental tradeoff between fitting the preference data and maintaining generative capabilities acquired during pre-training [22] and align with prior observations that adding on-policy preference data can make offline learning more effective [50, 65, 24, 53].
(a) Ranking accuracy and win rate of various Pythia 2.8B checkpoints during a single epoch of DPO training, versus the distance travelled by the model weights $\theta_{t}$ from the initialization.
(b) Ranking accuracy and win rate of various $\gamma$-scaled models (trained with $\mathcal{L}_{\text{DPO}}^{\gamma}$ for 5 epochs), versus the distance travelled by the model weights $\theta_{\gamma}$ from the initialization.
Figure 4: When the model weights have not travelled far from $\theta_{\text{Ref}}$, ranking accuracy and win rate increase together. $\theta_{t}$ represents the model weights at checkpoint $t$ during DPO training, and $\theta_{\gamma}$ represents the weights for a model trained to convergence with ${\mathcal{L}}_{\text{DPO}}^{\gamma}$.
@@ 6 Related Work
Analyses of Preference Learning Algorithms
Many works have investigated the role of the preference dataset [56, 61], the reliability of the evaluations [69, 29], and the confounding factor of response length [45, 10, 57, 37]. Theoretical works have unified the many preference learning algorithms into clear taxonomies that permit analysis and, sometimes, yield new variants [2, 61, 48, 50, 63, 34]. Several works study idiosyncrasies of preference learning, such as why DPO decreases the likelihood of both rejected and chosen outputs from the dataset [42, 13, 35] and why RLHF exhibits vanishing gradients [44]. In contrast, our work approaches understanding DPO and RLHF through the lens of ranking accuracy, and our findings emphasize the role of the reference model regularization in preference learning. Relatedly, SliC-HF [67], CPO [59], and pairwise cringe loss [60] optimize log probability margins $\log\pi(y_{w}|x)-\log\pi(y_{l}|x)$, effectively removing the regularization toward the reference model. Liu et al. [32] recommend using the reference model at inference time to exert more granular control over the regularization. Meng et al. [34] remove the $\pi_{\text{Ref}}$ terms altogether and use a target reward margin to prevent over-optimization. Additionally, Chennakesavalu et al. [6] design a DPO-like objective that includes an additional hyperparameter controlling the strength of the $\pi_{\text{Ref}}$ terms, similar to our $\mathcal{L}_{DPO}^{\gamma}$ objective (Eq. 4). Tang et al. [50] also analyze the role of regularizing toward a reference model, though our work focuses the effect of this regularization on ranking accuracy.
On-policy and Off-policy Preference Learning
@@ 7 Discussion
Our work highlights the significant but nuanced relationship between preference learning and ranking accuracy. We have demonstrated both theoretically and empirically that RLHF and DPO struggle to teach the model to correctly rank preferred and dispreferred outputs, even in the training dataset. Although the learning objective promotes high ranking accuracy in theory (Proposition 2.6), we observed a prominent alignment gap resulting from the poor conditioning of reference models. We then drew connections between the off-policy nature of ranking accuracy and the on-policy evaluations of win rate, identifying specific scenarios in which on-policy behavior can or cannot be reliably predicted by off-policy behavior. App. 8 details the limitations of our work.
@@ 9 Acknowledgements
We thank Sanjeev Arora, Tianyu Gao, Eric Mitchell, Richard Pang, and Mengzhou Xia for helpful discussions during the development of this project. We also thank Nikita Nangia, Eshaan Nichani, and Alexander Wettig for their help in proofreading the work. This work was supported by National Science Foundation Award 1922658, the Samsung Advanced Institute of Technology (under the project Next Generation Deep Learning: From Pattern Recognition to AI), NIH/NHLBI Award R01HL148248, NSF CAREER Award 2145542, ONR N00014-23-1-2634, Apple, and Google. SM is additionally supported by ONR, NSF, and DARPA. We also thank Princeton Language and Intelligence (PLI) for computing resources and OpenAI credits, and NYU High Performance Computing (HPC) for computing resources and in-kind support.